### v1

# references.bib

Abstract.
In this work, we develop a new compatible finite element formulation of the thermal shallow water equations that conserves energy and mathematical entropies given by buoyancy–related quadratic tracer variances. Our approach relies on restating the governing equations to enable discontinuous approximations of thermodynamic variables and a variational continuous time integration. A key novelty is the inclusion of centred and upwinded fluxes. The proposed semi-discrete system conserves discrete entropy for centred fluxes, monotonically damps entropy for upwinded fluxes, and conserves energy. The fully discrete scheme reflects entropy conservation at the continuous level. The ability of a new linearised Jacobian, which accounts for both centred and upwinded fluxes, to capture large variations in buoyancy and simulate thermally unstable flows for long periods of time is demonstrated for two different transient case studies. The first involves a thermogeostrophic instability where including upwinded fluxes is shown to suppress spurious oscillations while successfully conserving energy and monotonically damping entropy. The second is a double vortex where a constrained fully discrete formulation is shown to achieve exact entropy conservation in time.
Key words and phrases:
thermal shallow water equations, compatible finite elements, entropy conservation, Poisson systems, Casimir conservation
Corresponding author
1. Introduction
Mathematical entropies, or entropy, are convex functionals arising from a positive definite Hessian [Fisher2013, Leveque1992, Chapter 3] that correspond to quadratic buoyancy–related tracer invariants for the thermal shallow water equations [Ricardo2023, Ricardo2024-dg]. The thermal shallow water equations are a useful stepping stone from simpler atmospheric systems, like the rotating shallow water equations, to the full three-dimensional compressible Euler equations typically used to describe atmospheric dynamics in operational weather models [Maynard2020, Adams2019, Lee2021]. The thermal shallow equations are analogous to the compressible Euler equations with an identical Poisson bracket and entropy, and include a thermodynamic scalar quantity reflective of temperature [Eldred2019, Ricardo2023, Ricardo2024-dg]. Such thermodynamic quantities are not included in the rotating shallow water equations [Cotter2014, BauerCotter2018, Mcrae2013, Wimmer2020], meaning these other models do not capture the effect of thermodynamic transport on the horizontal pressure gradient [Kurganov2021]. In this study, we develop a novel finite element approximation of the thermal shallow water equations that provably conserves energy and entropy at the semi-discrete level. Large scale simulations show the resulting scheme is conservative, captures turbulent dynamics, and can stably simulate non-linear flow over long time scales where a mature turbulent state is reached.
[…]
Entropy conservation requires preserving certain conformity requirements in space and time that may not be inherited by the numerical approximations. A previous entropy conserving study uses a mixed finite element method in conjunction with Galerkin projections to enforce the necessary regularity of buoyancy fluxes [Ricardo2023]. Such continuity requirements are achieved in a different study via a discontinuous Galerkin method [Ricardo2024-dg]. In the current study, we consider a compatible finite element method that includes internal element boundary fluxes, cast in either a centred or upwinded form. The centred fluxes conserve energy and entropy, while the upwinded fluxes are proven to conserve energy and monotonically damp entropy. Care is taken to derive buoyancy terms that ensure entropy exchanges are balanced in both space and time for all forcing terms. In doing so, we ensure the only source of entropy conservation error is from temporal derivatives. The semi-discrete formulation also conserves total mass, buoyancy, vorticity, and supports compatible advection of buoyancy, thus satisfying the criteria of a compatible finite element discretisation for the thermal shallow water equations [Eldred2019, Ricardo2024-dg].
[…]
This article is structured as follows: in Section 2 we reformulate the thermal shallow water equations at the continuous level and analyse the regularity requirements for continuous entropy conservation. Finite element approximations are derived in Sections 3 and 4, where conservation at the semi- and fully discrete levels is evaluated. In Section 5, convergence of the new scheme under $h$-$p$ refinement is demonstrated using a steady thermogeostrophic balance test case [Eldred2019]. By considering small perturbations from the solution at progressive time levels instead of the usual mean flow state [Eldred2019, Lee2022], we propose a new linearised Jacobian and quasi-Newton approach that shows robust convergence in the presence of large variations in buoyancy for both centred and upwinded fluxes. Thus, we stably simulate thermally unstable flows to well evolved turbulent states, as demonstrated for transient case studies involving a thermogeostrophic instability [Eldred2019, Gouzien2017, Kurganov2021, Zeitlin2018] and a double vortex [Giorgetta2009, Eldred2019]. Conservation of invariants over long time scales is assessed, and Lagrange multipliers are used to correct small temporal losses in entropy conservation for centred numerical fluxes. The inclusion of upwinded numerical fluxes is shown to suppress spurious oscillations while conserving energy and monotonically damping entropy.
2. Thermal shallow water equations
In this section, we introduce the thermal shallow water equations and review their conservation properties. The thermal shallow water equations describe the evolution of fluid velocity $\boldsymbol{u}(\boldsymbol{x},t)$ and depth $\varphi(\boldsymbol{x},t)$ with respect to buoyancy transport. There are two types of buoyancy transport [Eldred2019, Ricardo2023, Ricardo2024-dg]. The first is material transport of buoyancy $b(\boldsymbol{x},t)=g\rho(\boldsymbol{x},t)/\bar{\rho}$ as a function of density $\rho(\boldsymbol{x},t)$, vertically averaged density $\bar{\rho}$, and gravity $g$ [Zeitlin2018]. The second is flux transport of density weighted buoyancy $B(\boldsymbol{x},t)=\varphi(\boldsymbol{x},t)b(\boldsymbol{x},t)$. Both support non-canonical Hamiltonian formulations and reduce to the rotating shallow water equations when $b(\boldsymbol{x},t)=g$. We consider flux transport of $B(\boldsymbol{x},t)$. This yields the vector invariant form of the thermal shallow water equations for which the total energy, or Hamiltonian, is
$\mathcal{H}(t)$ $=\int_{\Omega}\left(\frac{1}{2}\varphi(\boldsymbol{x},t)\boldsymbol{u}(\boldsymbol{x},t)\cdot\boldsymbol{u}(\boldsymbol{x},t)+\frac{1}{2}\varphi(\boldsymbol{x},t)B(\boldsymbol{x},t)\right).$ (1)
2.1. Continuous system
[…]
$\frac{\partial\boldsymbol{u}(\boldsymbol{x},t)}{\partial t}+q(\boldsymbol{x},t)\boldsymbol{F}(\boldsymbol{x},t)^{\perp}+\nabla\Phi(\boldsymbol{x},t)+b(\boldsymbol{x},t)\nabla\vartheta(\boldsymbol{x},t)$ $=0,\qquad\text{in }\Omega\times J,$ (2a)
$\frac{\partial\varphi(\boldsymbol{x},t)}{\partial t}+\nabla\cdot\boldsymbol{F}(\boldsymbol{x},t)$ $=0,\qquad\text{in }\Omega\times J,$ (2b)
$\frac{\partial B(\boldsymbol{x},t)}{\partial t}+\nabla\cdot\left(b(\boldsymbol{x},t)\boldsymbol{F}(\boldsymbol{x},t)\right)$ $=0,\qquad\text{in }\Omega\times J,$ (2c)
where $q(\boldsymbol{x},t)$ is the potential vorticity. The mass flux $\boldsymbol{F}(\boldsymbol{x},t)$, Bernoulli potential $\Phi(\boldsymbol{x},t)$, and temperature $\vartheta(\boldsymbol{x},t)$, are the variational derivatives of $\mathcal{H}(t)$ with respect to $\boldsymbol{u}(\boldsymbol{x},t)$, $\varphi(\boldsymbol{x},t)$, $B(\boldsymbol{x},t)$, respectively. The prognostic variables, $\boldsymbol{u}(\boldsymbol{x},t),\varphi(\boldsymbol{x},t),B(\boldsymbol{x},t)$, solve the time dependent problem 2. The diagnostic variables, $\boldsymbol{F}(\boldsymbol{x},t),\Phi(\boldsymbol{x},t),\vartheta(\boldsymbol{x},t),q(\boldsymbol{x},t),b(\boldsymbol{x},t)$, solve the algebraic constraints
$\boldsymbol{F}(\boldsymbol{x},t)$ $=\varphi(\boldsymbol{x},t)\boldsymbol{u}(\boldsymbol{x},t),$ $\Phi(\boldsymbol{x},t)$ $=\frac{1}{2}\boldsymbol{u}(\boldsymbol{x},t)\cdot\boldsymbol{u}(\boldsymbol{x},t)+\frac{1}{2}B(\boldsymbol{x},t),$ $\vartheta(\boldsymbol{x},t)$ $=\frac{1}{2}\varphi(\boldsymbol{x},t),$
$q(\boldsymbol{x},t)$ $=\frac{\nabla^{\perp}\cdot\boldsymbol{u}(\boldsymbol{x},t)+f}{\varphi(\boldsymbol{x},t)},$ $b(\boldsymbol{x},t)$ $=\frac{B(\boldsymbol{x},t)}{\varphi(\boldsymbol{x},t)},$ (3)
where $f$ is the Coriolis parameter.
2.2. System invariants
Using a skew symmetric operator to express 2 in terms of $\mathcal{H}(t)$ means energy is conserved [Eldred2019, Ricardo2023, Ricardo2024-dg, Wimmer2020, BauerCotter2018]. Skew symmetric operators, which represent non-canonical Poisson brackets for compressible fluids, have a nullspace that contains the variational derivatives of additional system invariants, known as Casimirs [Shepherd1990, Salmon1998]. These include total mass $\mathcal{M}(t)=\int_{\Omega}\varphi(\boldsymbol{x},t)$, buoyancy $\mathcal{B}(t)=\int_{\Omega}B(\boldsymbol{x},t)$, vorticity $\mathcal{V}(t)=\int_{\Omega}\varphi(\boldsymbol{x},t)q(\boldsymbol{x},t)$ and higher order moments of buoyancy such as entropy
$\mathcal{S}(t)$ $=\int_{\Omega}\frac{1}{2}b(\boldsymbol{x},t)^{2}\varphi(\boldsymbol{x},t).$ (4)
To motivate the discrete approximation proposed in Sections 3 and 4 below, we discuss entropy conservation at the continuous level. Hereafter, we drop the dependence on space and time. Assuming $b,\varphi,B\in H^{1}(J)$, the total time derivative of $\mathcal{S}$ is
$\frac{\mathrm{d}\mathcal{S}}{\mathrm{d}t}=\int_{\Omega}b\varphi\frac{\partial b}{\partial t}+\int_{\Omega}\frac{1}{2}b^{2}\frac{\partial\varphi}{\partial t}=\int_{\Omega}\left(b\frac{\partial B}{\partial t}-b^{2}\frac{\partial\varphi}{\partial t}\right)+\int_{\Omega}\frac{1}{2}b^{2}\frac{\partial\varphi}{\partial t}$ $=\int_{\Omega}\frac{\delta\mathcal{S}}{\delta B}\frac{\partial B}{\partial t}+\int_{\Omega}\frac{\delta\mathcal{S}}{\delta\varphi}\frac{\partial\varphi}{\partial t},$ (5)
where $\delta\mathcal{S}/\delta B=b$ and $\delta\mathcal{S}/\delta\varphi=-b^{2}/2$ are the variational derivatives of entropy. Next, substitute 2b and 2c into 5. Expanding spatial derivatives and assuming $\nabla(b^{2})\in H^{1}(\Omega)$, $\boldsymbol{F}\in H(\mathrm{div},\Omega)$ in order to apply integration by parts gives
$\frac{\mathrm{d}\mathcal{S}}{\mathrm{d}t}$ $=\int_{\Omega}\frac{1}{2}b^{2}\nabla\cdot\boldsymbol{F}-\int_{\Omega}b\nabla\cdot(b\boldsymbol{F})=\int_{\Omega}\frac{1}{2}\nabla(b^{2})\cdot\boldsymbol{F}-\int_{\Omega}b\boldsymbol{F}\cdot\nabla b=0.$ (6)
The above analysis is used to determine the regularity requirements for discrete entropy conservation. For 5 to hold discretely in time, we consider a temporally continuous approximation of $b$ and the prognostic variables. Thus, entropy is continuous cubic polynomial in time, which is not conserved in general at the discrete level by the chosen time integrator [Cohen2011]. The spatial approximation of prognostic variables is determined by a discrete de Rham complex [Cotter2023], where $\nabla\cdot\boldsymbol{u},\varphi,B$ are spatially discontinuous. For the vector invariant form of the thermal shallow water equations, $b$ is typically also spatially discontinuous [Ricardo2023, Eldred2019]. This choice means entropy conservation in 6 does not hold at the semi-discrete level [Ricardo2023]. While the material form of the thermal shallow water equations facilitates spatially continuous approximations of $b$, the associated variational derivative of the Hamiltonian is not collocated with the pressure gradient, and energy is not readily conserved [Eldred2019].
To allow for spatially discontinuous approximations of $b$, we define a numerical approximation of 2 such that integration by parts is not required to prove semi-discrete entropy conservation. Using the chain rule to reformulate the buoyancy equation 2c and reciprocating modifications in the momentum equation 2a, we restate 2 as [Ricardo2023]
$\frac{\partial\boldsymbol{u}}{\partial t}+q\boldsymbol{F}^{\perp}+\nabla\Phi+\frac{1}{2}b\nabla\vartheta+\frac{1}{2}\nabla(b\vartheta)-\frac{1}{2}\vartheta\nabla b$ $=0,$ (7a)
$\frac{\partial\varphi}{\partial t}+\nabla\cdot\boldsymbol{F}$ $=0,$ (7b)
$\frac{\partial B}{\partial t}+\frac{1}{2}\nabla\cdot(b\boldsymbol{F})+\frac{1}{2}b\nabla\cdot\boldsymbol{F}+\frac{1}{2}\boldsymbol{F}\cdot\nabla b$ $=0.$ (7c)
Skew symmetry is maintained in 7 so that energy is conserved, assuming periodic spatial boundary conditions. Substituting 7b and 7c into 5, yields entropy conservation at the continuous level as in 6. At the discrete level, we apply integration by parts to $\nabla\cdot(b\boldsymbol{F})$ in 7c and account for jumps across interior element boundaries such that the resulting term readily cancels with the jumps related to the $\boldsymbol{F}\cdot\nabla b$ term in 7c. The $b\nabla\cdot\boldsymbol{F}$ term in 7c is div-conforming at the discrete level, and does not contribute to interior element boundaries. To ensure discrete energy and entropy conservation, we design skew symmetric jump terms by exploiting the regularity of $b$ at the continuous level.
3. Semi-discrete formulation
We now present a compatible finite element formulation of the thermal shallow water equations 7, and prove semi-discrete energy and entropy conservation. Skew symmetric upwinded numerical fluxes, which provably conserve energy and monotonically damp entropy, are proposed to smooth spurious oscillations that can occur as simulations progress to mature turbulent states.
3.1. Notation
[…]
$\llbracket x\rrbracket$ $=x^{+}\boldsymbol{n}^{+}+x^{-}\boldsymbol{n}^{-},$$\left\{x\right\}$ $=\frac{1}{2}\left(x^{+}+x^{-}\right)\text{ on }e\in\mathcal{E}_{0},$
$\llbracket\boldsymbol{y}\rrbracket$ $=\boldsymbol{y}^{+}\boldsymbol{n}^{+}+\boldsymbol{y}^{-}\boldsymbol{n}^{-},$$\left\{\boldsymbol{y}\right\}$ $=\frac{1}{2}\left(\boldsymbol{y}^{+}+\boldsymbol{y}^{-}\right)\text{ on }e\in\mathcal{E}_{0},$
where $\mathcal{E}_{0}$ is the set of interior edges, $x$ is a scalar valued function, $\boldsymbol{y}$ is a vector valued function, $\cdot^{\pm}$ is the restriction to $K^{\pm}$, and $\boldsymbol{n}^{+}=-\boldsymbol{n}^{-}$ is the outward pointing spatial normal vectors of $K^{\pm}$ on $e$.
3.2. Finite element spaces
[…]
We consider quadrilateral meshes where $\mathbb{V}_{0}=\mathcal{P}_{p+1}(\mathcal{T}_{h}^{x})$ is the scalar continuous Lagrangian finite element space of piecewise polynomials of order $p+1$, $\mathbb{V}_{1}=\mathcal{RT}_{p}(\mathcal{T}_{h}^{x})$ is the polynomial Raviart-Thomas space of order $p$ [RT1977], and $\mathbb{V}_{2}=\mathcal{P}_{p}^{-}(\mathcal{T}_{h}^{x})$ is the space of piecewise discontinuous polynomials of order $p$. The discussion hereafter applies to any other choice of compatible finite element spaces for quadrilateral and triangular meshes [Arnold2018].
3.3. Finite element approximation
The semi-discrete finite element formulation of the thermal shallow water equations in 7 is: find $\boldsymbol{u}_{h}\in\mathbb{V}_{1}$, $\varphi_{h},B_{h}\in\mathbb{V}_{2}$, such that
$\left(\frac{\partial\boldsymbol{u}_{h}}{\partial t},\psi_{\boldsymbol{u}_{h}}\right)$ $+\left(q_{h},\boldsymbol{F}_{h}^{\perp}\cdot\psi_{\boldsymbol{u}_{h}}\right)-\left(\nabla\cdot\psi_{\boldsymbol{u}_{h}},\Phi_{h}\right)$
$-g(\psi_{\boldsymbol{u}_{h}},b_{h},\widetilde{b_{h}},\vartheta_{h})-s(\psi_{\boldsymbol{u}_{h}},b_{h},\vartheta_{h})=0,$ $\forall\psi_{\boldsymbol{u}_{h}}\in\mathbb{V}_{1},$ (8a)
$\left(\frac{\partial\varphi_{h}}{\partial t},\phi_{\varphi_{h}}\right)$ $+\left(\nabla\cdot\boldsymbol{F}_{h},\phi_{\varphi_{h}}\right)=0,$ $\forall\phi_{\varphi_{h}}\in\mathbb{V}_{2},$ (8b)
$\left(\frac{\partial B_{h}}{\partial t},\phi_{B_{h}}\right)$ $+g(\boldsymbol{F}_{h},b_{h},\widetilde{b_{h}},\phi_{B_{h}})+s(\boldsymbol{F}_{h},b_{h},\phi_{B_{h}})=0,$ $\forall\phi_{B_{h}}\in\mathbb{V}_{2},$ (8c)
together with periodic spatial boundary conditions. The forms $g(\cdot,b_{h},\widetilde{b_{h}},\cdot)$ and $s(\cdot,b_{h},\cdot)=s_{\mathrm{c}}(\cdot,b_{h},\cdot)+s_{\mathrm{up}}(\cdot,b_{h},\cdot)$ are the discretisation of the advection and stabilisation terms, which consists of a centred and upwinded flux, as follows:
$g(\boldsymbol{w}_{h},b_{h},\widetilde{b_{h}},\phi_{h})$ $=-\frac{1}{2}\left(b_{h},\boldsymbol{w}_{h}\cdot\nabla_{h}\phi_{h}\right)+\frac{1}{2}\left(\widetilde{b_{h}}\phi_{h},\nabla\cdot\boldsymbol{w}_{h}\right)+\frac{1}{2}\left(\phi_{h},\nabla_{h}b_{h}\cdot\boldsymbol{w}_{h}\right),$ (9a)
$s_{\mathrm{c}}(\boldsymbol{w}_{h},b_{h},\phi_{h})$ $=\frac{1}{2}\left\langle\left\{\boldsymbol{w}_{h}b_{h}\right\},\llbracket\phi_{h}\rrbracket\right\rangle-\frac{1}{2}\left\langle\left\{\boldsymbol{w}_{h}\phi_{h}\right\},\llbracket b_{h}\rrbracket\right\rangle,$ (9b)
$s_{\mathrm{up}}(\boldsymbol{w}_{h},b_{h},\phi_{h})$ $=\frac{1}{2}\left\langle\alpha(\boldsymbol{w}_{h})\llbracket\phi_{h}\rrbracket,\llbracket b_{h}\rrbracket\right\rangle,$ (9c)
for all $\boldsymbol{w}_{h}\in\mathbb{V}_{1}$, $\phi_{h}\in\mathbb{V}_{2}$, where $\widetilde{b_{h}}$ is a specific formulation of the buoyancy introduced in 25 below, and $\alpha(\boldsymbol{w}_{h})\coloneqq|\boldsymbol{w}_{h}\cdot\boldsymbol{n}^{+}|/2$ is the upwinding parameter [Brezzi2004]. Centred fluxes correspond to $\alpha=0$ and upwinded fluxes correspond to $\alpha>0$. The upwinded fluxes in 9c align with other studies that consider hyperbolic systems involving advection [Brezzi2004, Eldred2019], and maintain skew symmetry in 8. Upwinding is not required in the continuity equation since the corresponding linearised equation is balanced with respect to left and right solutions [Vallis2006], and $\boldsymbol{F}_{h}\in\mathbb{V}_{1}$ is continuous across element boundaries.
The semi-discrete approximation of the diagnostic variables in 3 is: find $\boldsymbol{F}_{h}\in\mathbb{V}_{1}$, $\Phi_{h},\vartheta_{h},b_{h}\in\mathbb{V}_{2}$, $q_{h}\in\mathbb{V}_{0}$, such that
$\left(\boldsymbol{F}_{h},\psi_{\boldsymbol{F}_{h}}\right)$ $=\left(\varphi_{h}\boldsymbol{u}_{h},\psi_{\boldsymbol{F}_{h}}\right),$ $\forall\psi_{\boldsymbol{F}_{h}}\in\mathbb{V}_{1},$ (10a)
$\left(\Phi_{h},\phi_{\Phi_{h}}\right)$ $=\left(\frac{1}{2}\boldsymbol{u}_{h}\cdot\boldsymbol{u}_{h}+B_{h},\phi_{\Phi_{h}}\right),$ $\forall\phi_{\Phi_{h}}\in\mathbb{V}_{2},$ (10b)
$\left(\vartheta_{h},\phi_{\vartheta_{h}}\right)$ $=\left(\frac{1}{2}\varphi_{h},\phi_{\vartheta_{h}}\right),$ $\forall\phi_{\vartheta_{h}}\in\mathbb{V}_{2},$ (10c)
$\left(b_{h}\varphi_{h},\phi_{b_{h}}\right)$ $=\left(B_{h},\phi_{b_{h}}\right),$ $\forall\phi_{b_{h}}\in\mathbb{V}_{2},$ (10d)
$\left(q_{h}\varphi_{h},\xi_{q_{h}}\right)$ $=-\left(\nabla^{\perp}\xi_{q_{h}},\boldsymbol{u}_{h}\right)+\left(f,\xi_{q_{h}}\right),$ $\forall\xi_{q_{h}}\in\mathbb{V}_{0},$ (10e)
where 10c holds point-wise and is not required in practice. The final diagnostic is: find $\widetilde{b_{h}}\in\mathbb{V}_{2}$, such that
$\left(\widetilde{b_{h}}b_{h},\widetilde{\phi}_{h}\right)$ $=\left(b_{h}b_{h},\widetilde{\phi}_{h}\right),$ $\forall\widetilde{\phi}_{h}\in\mathbb{V}_{2},$ (11)
where the projection of the quadratic term $b_{h}b_{h}$ into $\mathbb{V}_{2}$ is accounted for since $\widetilde{\phi}_{h}\in\mathbb{V}_{2}$. While entropy can be conserved at the semi-discrete level by replacing $\widetilde{b_{h}}$ with $b_{h}$ [Ricardo2023, Ricardo2024-dg], we show in Section 4.3 that the additional diagnostic 11 ensures the forcing terms exactly conserve entropy in time. Thus, the only source of temporal conservation error is the approximation of time derivatives.
We now prove the semi-discrete formulation is a consistent finite element approximation of the thermal shallow water equations that supports compatible advection of buoyancy.
Proposition 3.1.
Solutions of the continuous system are consistent with the semi-discrete formulation.
Proof.
To show smooth solutions of the continuous system satisfy the semi-discrete system, discrete quantities in 8 are replaced by their continuous analogue. Equation 8 contains terms involving the time derivative of the prognostic variables, for which the consistency is trivial, and all other terms relate to the diagnostic variables. We now elaborate on the consistency of the forms $g(\cdot,b_{h},\widetilde{b_{h}},\cdot)$ and $s(\cdot,b_{h},\cdot)$.
First we observe that $\llbracket b\rrbracket=0$, $\widetilde{b}=b$, and $\llbracket\vartheta\rrbracket=0$ for smooth solutions at the continuous level. For the momentum equation 8a, the terms in question reduce to
$-g(\psi_{\boldsymbol{u}},b,b,\vartheta)-s(\psi_{\boldsymbol{u}},b,\vartheta)$ $=-g(\psi_{\boldsymbol{u}},b,b,\vartheta)=\left(b,\nabla\vartheta\cdot\psi_{\boldsymbol{u}}\right),$
after applying integration by parts and the chain rule. Consistency readily follows noting that $\nabla\cdot\psi_{\boldsymbol{u}}\Phi$ is div-conforming. We proceed analogously for the density weighted buoyancy equation 8c. ∎
Proposition 3.2.
The semi-discrete formulation supports compatible advection of buoyancy.
Proof.
[…]
Setting $b_{h}=1$ in 10d and 11 yields $\varphi_{h}=B_{h}$ point-wise and $\widetilde{b_{h}}=1$ respectively. Next taking $b_{h}=\widetilde{b_{h}}=1$ in 8c, and expanding $g(\boldsymbol{F}_{h},1,1,\phi_{B_{h}})$ and $s(\boldsymbol{F}_{h},1,\phi_{B_{h}})$ gives
$\left(\frac{\partial B_{h}}{\partial t},\phi_{B_{h}}\right)$ $+\frac{1}{2}\left(\phi_{h},\nabla\cdot\boldsymbol{F}_{h}\right)-\frac{1}{2}\left(1,\boldsymbol{F}_{h}\cdot\nabla_{h}\phi_{h}\right)+\frac{1}{2}\left\langle\left\{\boldsymbol{F}_{h}\right\},\llbracket\phi_{h}\rrbracket\right\rangle=0,$ $\forall\phi_{B_{h}}\in\mathbb{V}_{2},$ (12)
since $\llbracket 1\rrbracket=\nabla_{h}1=0$ and $\left\{1\right\}=1$. Applying integration by parts yields 8b since $\varphi_{h}=B_{h}$. ∎
3.4. Semi-discrete conservation
We now prove the main result of this study: the semi-discrete system 8 conserves discrete entropy when $\alpha=0$, monotonically damps entropy when $\alpha>0$, and conserves energy always.
Proposition 3.3.
The semi-discrete system conserves energy.
Proof.
Discrete energy at any time $t$ is
$\mathcal{H}_{h}(t)$ $=\int_{\Omega}\left(\frac{1}{2}\varphi_{h}\boldsymbol{u}_{h}\cdot\boldsymbol{u}_{h}+\frac{1}{2}\varphi_{h}B_{h}\right),\qquad\forall t\in J.$ (13)
Differentiating $\mathcal{H}_{h}(t)$ with respect to time gives
$\frac{\partial\mathcal{H}}{\partial t}$ $=\left(\frac{\delta\mathcal{H}_{h}}{\delta\boldsymbol{u}_{h}},\frac{\partial\boldsymbol{u}_{h}}{\partial t}\right)+\left(\frac{\delta\mathcal{H}_{h}}{\delta\varphi_{h}},\frac{\partial\varphi_{h}}{\partial t}\right)+\left(\frac{\delta\mathcal{H}_{h}}{\delta B_{h}},\frac{\partial B_{h}}{\partial t}\right),$ (14)
where
$\frac{\delta\mathcal{H}_{h}}{\delta\boldsymbol{u}_{h}}\in\mathbb{V}_{1}:$ $\left(\frac{\delta\mathcal{H}_{h}}{\delta\boldsymbol{u}_{h}},\phi_{\boldsymbol{u}_{h}}\right)$ $=\left(\varphi_{h}\boldsymbol{u}_{h},\phi_{\boldsymbol{u}_{h}}\right)=\left(\boldsymbol{F}_{h},\phi_{\boldsymbol{u}_{h}}\right),$ $\forall\phi_{\boldsymbol{u}_{h}}\in\mathbb{V}_{1},$ (15a)
$\frac{\delta\mathcal{H}_{h}}{\delta\varphi_{h}}\in\mathbb{V}_{2}:$ $\left(\frac{\delta\mathcal{H}_{h}}{\delta\varphi_{h}},\phi_{\varphi_{h}}\right)$ $=\left(\frac{1}{2}\boldsymbol{u}_{h}\cdot\boldsymbol{u}_{h}+B_{h},\phi_{\varphi_{h}}\right)=\left(\Phi_{h},\phi_{\varphi_{h}}\right),$ $\forall\phi_{\varphi_{h}}\in\mathbb{V}_{2},$ (15b)
$\frac{\delta\mathcal{H}_{h}}{\delta B_{h}}\in\mathbb{V}_{2}:$ $\left(\frac{\delta\mathcal{H}_{h}}{\delta B_{h}},\phi_{B_{h}}\right)$ $=\left(\frac{1}{2}\varphi_{h},\phi_{B_{h}}\right)=\left(\vartheta_{h},\phi_{B_{h}}\right),$ $\forall\phi_{B_{h}}\in\mathbb{V}_{2},$ (15c)
are the functional derivatives of $\mathcal{H}_{h}(t)$ that hold point-wise. Substituting 15 into 14 and evaluating the resulting expression using $\psi_{\boldsymbol{u}_{h}}=\boldsymbol{F}_{h}$, $\phi_{\varphi_{h}}=\Phi_{h}$, $\phi_{B_{h}}=\vartheta_{h}$ in 8 gives ${\partial\mathcal{H}_{h}}/{\partial t}=0$. ∎
Proposition 3.4.
The semi-discrete system conserves entropy when $\alpha=0$, and monotonically damps entropy otherwise for $\alpha>0$.
Proof.
Discrete entropy at any time $t$ is
$\mathcal{S}_{h}(t)$ $=\int_{\Omega}\frac{1}{2}b_{h}b_{h}\varphi_{h},\qquad\forall t\in J.$ (16)
Under the assumption of continuity of time, 5 gives the time derivative of $\mathcal{S}_{h}$ as
$\frac{\partial\mathcal{S}_{h}}{\partial t}$ $=\left(\frac{\delta\mathcal{S}_{h}}{\delta\varphi_{h}},\frac{\partial\varphi_{h}}{\partial t}\right)+\left(\frac{\delta\mathcal{S}_{h}}{\delta{B_{h}}},\frac{\partial B_{h}}{\partial t}\right).$ (17)
In contrast to energy, entropy involves both prognostic and diagnostic variables. Evaluating the discrete functional derivatives requires several steps as follows. First differentiating 16 with respect to $\varphi_{h}$ and $b_{h}$ gives
$\left(\frac{\partial\mathcal{S}_{h}}{\partial\varphi_{h}},\phi_{\varphi_{h}}\right)$ $=\left(\frac{1}{2}b_{h}b_{h},\phi_{\varphi_{h}}\right),$ $\forall\phi_{\varphi_{h}}\in\mathbb{V}_{2},$ (18a)
$\left(\frac{\partial\mathcal{S}_{h}}{\partial b_{h}},\phi_{b_{h}}\right)$ $=\left(b_{h}\varphi_{h},\phi_{b_{h}}\right),$ $\forall\phi_{b_{h}}\in\mathbb{V}_{2}.$ (18b)
Further differentiating 10d with respect to $\varphi_{h}$ and $B_{h}$ yields
$\left(\frac{\partial b_{h}}{\partial\varphi_{h}}\varphi_{h},\phi_{b_{h}}\right)$ $=\left(-b_{h}\phi_{\varphi_{h}},\phi_{b_{h}}\right),$ $\forall\phi_{b_{h}},\phi_{\varphi_{h}}\in\mathbb{V}_{2},$ (19a)
$\left(\frac{\partial b_{h}}{\partial B_{h}}\varphi_{h},\phi_{b_{h}}\right)$ $=\left(\phi_{B_{h}},\phi_{b_{h}}\right),$ $\forall\phi_{b_{h}},\phi_{B_{h}}\in\mathbb{V}_{2}.$ (19b)
Thus, the variational derivatives of $\mathcal{S}_{h}(t)$ are:
Find $\text{Find }\frac{\delta\mathcal{S}_{h}}{\delta\varphi_{h}}\in\mathbb{V}_{2}:\left(\frac{\delta\mathcal{S}_{h}}{\delta\varphi_{h}},\phi_{\varphi_{h}}\right)$ $=\left(\frac{\partial\mathcal{S}_{h}}{\partial b_{h}},\frac{\partial b_{h}}{\partial\varphi_{h}}\right)+\left(\frac{\partial\mathcal{S}_{h}}{\partial\varphi_{h}},\phi_{\varphi_{h}}\right)=\left(-\frac{1}{2}b_{h}b_{h},\phi_{\varphi_{h}}\right),~{}\forall\phi_{\varphi_{h}}\in\mathbb{V}_{2},$ (20a)
Find $\text{Find }\frac{\delta\mathcal{S}_{h}}{\delta B_{h}}\in\mathbb{V}_{2}:\left(\frac{\delta\mathcal{S}_{h}}{\delta B_{h}},\phi_{B_{h}}\right)$ $=\left(\frac{\partial\mathcal{S}_{h}}{\partial b_{h}},\frac{\partial b_{h}}{\partial B_{h}}\right)=\left(b_{h},\phi_{B_{h}}\right),~{}~{}~{}~{}\forall\phi_{B_{h}}\in\mathbb{V}_{2}.$ (20b)
In obtaining 20a, we set $\phi_{b_{h}}={\partial b_{h}}/{\partial\varphi_{h}}$ in 18b, $\phi_{\varphi_{h}}=b_{h}$ in 19a, and combine the result with 18a. Deriving 20b requires taking $\phi_{b_{h}}={\partial b_{h}}/{\partial B_{h}}$ in 18b and $\phi_{b_{h}}=b_{h}$ in 19b.
To evaluate 17, set $\phi_{\varphi_{h}}=-\Pi_{\mathbb{V}_{2}}(b_{h}b_{h})/2$ in 8b and $\phi_{B_{h}}=b_{h}$ in 8c, where $\Pi_{\mathbb{V}_{2}}$ is the projection into $\mathbb{V}_{2}$. Expanding $g(\boldsymbol{F}_{h},b_{h},\widetilde{b_{h}},b_{h})$ and $s(\boldsymbol{F}_{h},b_{h},b_{h})$ in the resulting expression shows integration by parts is not required to cancel terms. Thus,
$\frac{\partial\mathcal{S}_{h}}{\partial t}$ $=\frac{1}{2}\left(b_{h}b_{h},\nabla\cdot\boldsymbol{F}_{h}\right)-\frac{1}{2}\left(\widetilde{b_{h}}{b_{h}},\nabla\cdot\boldsymbol{F}_{h}\right)-s_{\mathrm{up}}(\boldsymbol{F}_{h},{b_{h}},{b_{h}}),$ (21)
where the projection into $\mathbb{V}_{2}$ cancels since $\nabla\cdot\boldsymbol{F}_{h}\in\mathbb{V}_{2}$. Using $\widetilde{\phi}_{h}=\nabla\cdot\boldsymbol{F}_{h}$ in 11 yields ${\partial\mathcal{S}_{h}}/{\partial t}=0$ when $\alpha=0$, and ${\partial\mathcal{S}_{h}}/{\partial t}<0$ when $\alpha>0$ since $s_{\mathrm{up}}(\boldsymbol{F}_{h},{b_{h}},{b_{h}})>0$ for $\alpha>0$. ∎
[…]
Discrete mass $\mathcal{M}_{h}(t)=\int_{\Omega}\varphi_{h}$ is conserved since setting $\phi_{\varphi_{h}}=1$ in 8b results in a flux form conservation law. Taking $\phi_{B_{h}}=1$ in 8c and applying integration by parts reveals the discrete total buoyancy, $\mathcal{B}_{h}=\int_{\Omega}B_{h}$, is only conserved when $\widetilde{b_{h}}=b_{h}$ point-wise. Similar to other finite element approximations of the thermal shallow water equations [Eldred2019], discrete vorticity, $\mathcal{V}_{h}=\int_{\Omega}\varphi_{h}q_{h}$, is conserved when $b_{h}$ is a constant. Thus, the semi-discrete formulation satisfies the criteria of a compatible spatial discretisation for the thermal shallow water equations [Eldred2019, Ricardo2024-dg].
4. Fully discrete formulation
In this section, we integrate the semi-discrete system 8 in time to develop a fully discrete formulation of the thermal shallow water equations. Temporal conservation is also analysed.
4.1. Poisson integrator
The semi-discrete system 8 is a Poisson system of the form
$\frac{\partial\boldsymbol{z}_{h}}{\partial t}=\mathbb{A}(\boldsymbol{z}_{h})\frac{\delta\mathcal{H}_{h}(\boldsymbol{z}_{h})}{\delta\boldsymbol{z}_{h}},$ (22)
[…]
Let $\mathcal{T}_{h}^{t}$ be a conforming partition of the temporal domain into $N$ discrete intervals, $J^{n}\coloneqq(t^{n-1},t^{n}]$. For simplicity, we assume all time intervals have equal length $\tau\coloneqq T/N$. The discrete time nodes are $t^{n}\coloneqq n\tau$ for $n=1,2,\ldots,N$. For linear approximations in time, the Poisson integrator [Cohen2011] is
$\frac{\boldsymbol{z}_{h}^{n}-\boldsymbol{z}_{h}^{n-1}}{\tau}=\mathbb{A}\left(\frac{\boldsymbol{z}_{h}^{n}+\boldsymbol{z}_{h}^{n-1}}{2}\right)\int_{0}^{1}\frac{\delta\mathcal{H}_{h}}{\delta\boldsymbol{z}_{h}}\left(\boldsymbol{z}_{h}^{n-1}+s\left(\boldsymbol{z}_{h}^{n}-\boldsymbol{z}_{h}^{n-1}\right)\right)\mathrm{d}s,$ (23)
where the discrete variational derivatives of the Hamiltonian are exactly integrated in time [BauerCotter2018, Eldred2019]. By construction of the Poisson integrator [Cohen2011], applying 23 to the semi-discrete system 8 yields temporal energy conservation. Unfortunately the chosen Poisson integrator [Cohen2011] does not generalise to cubic Casimirs, such as discrete entropy, and we are not aware of any time integrator that preserves higher order Casimirs in general for non-canonical Hamiltonian systems.
4.2. Temporal discretisation
[…]
Without loss of generality, let $J^{n}\in(0,1]$. We find temporal nodal values of $b_{h}^{n}\in\mathbb{V}_{2}$ using $\varphi_{h}^{n},B_{h}^{n}\in\mathbb{V}_{2}$ in the semi-discrete relation 10d. Then, $b_{h}^{n-1/2}=(b_{h}^{n-1}+b_{h}^{n})/2$ is the mean value of a linear polynomial in time for $t\in J^{n}$.
Thus, the fully discrete system is: find $\boldsymbol{u}_{h}^{n}\in\mathbb{V}_{1}$, $\varphi_{h}^{n},B_{h}^{n}\in\mathbb{V}_{2}$, such that
$\left(\boldsymbol{u}_{h}^{n},\psi_{\boldsymbol{u}_{h}}\right)-\left(\boldsymbol{u}_{h}^{n-1},\psi_{\boldsymbol{u}_{h}}\right)$ $+\tau\left(q_{h}^{n-1/2},\boldsymbol{F}_{h}^{\perp,n}\cdot\psi_{\boldsymbol{u}_{h}}\right)-\tau\left(\nabla\cdot\psi_{\boldsymbol{u}_{h}},\Phi_{h}^{n}\right)$
$-\tau g(\psi_{\boldsymbol{u}_{h}},b_{h}^{n-1/2},\widetilde{b_{h}^{n}},\vartheta_{h}^{n})-\tau s(\psi_{\boldsymbol{u}_{h}},b_{h}^{n-1/2},\vartheta_{h}^{n})=0,$ $\forall\psi_{\boldsymbol{u}_{h}}\in\mathbb{V}_{1},$ (24a)
$\left(\varphi_{h}^{n},\phi_{\varphi_{h}}\right)-\left(\varphi_{h}^{n-1},\phi_{\varphi_{h}}\right)$ $+\tau\left(\nabla\cdot\boldsymbol{F}_{h}^{n},\phi_{\varphi_{h}}\right)=0,$ $\forall\phi_{\varphi_{h}}\in\mathbb{V}_{2},$ (24b)
$\left(B_{h}^{n},\phi_{B_{h}}\right)-\left(B_{h}^{n-1},\phi_{B_{h}}\right)$ $+\tau g(\boldsymbol{F}_{h}^{n},b_{h}^{n-1/2},\widetilde{b_{h}^{n}},\phi_{B_{h}})+\tau s(\boldsymbol{F}_{h}^{n},b_{h}^{n-1/2},\phi_{B_{h}})=0,$ $\forall\phi_{B_{h}}\in\mathbb{V}_{2},$ (24c)
where $\widetilde{b_{h}^{n}}\in\mathbb{V}_{2}$ satisfies,
$\left(\widetilde{b_{h}^{n}}b_{h}^{n-1/2},\widetilde{\phi}_{h}\right)$ $=\left([b_{h}b_{h}]^{n-1/2},\widetilde{\phi}_{h}\right),$ $\forall\widetilde{\phi}_{h}\in\mathbb{V}_{2},$ (25)
and $[b_{h}b_{h}]^{n-1/2}=(b_{h}^{n-1}b_{h}^{n-1}+b_{h}^{n}b_{h}^{n})/2$ is the mean value of the square of $b_{h}$. Quadratures up to the degree required are used to evaluate $\boldsymbol{F}^{n}_{h}\in\mathbb{V}_{1}$, $\Phi^{n}_{h},\vartheta^{n}_{h}\in\mathbb{V}_{2}$, $q_{h}^{n-1/2}\in\mathbb{V}_{0}$ exactly in time using $\boldsymbol{u}_{h}^{n},\varphi_{h}^{n},B_{h}^{n}$ in 10 [Eldred2019, BauerCotter2018].
4.3. Temporal conservation
Discrete energy is conserved point-wise in time by construction of the Poisson integrator [Cohen2011]. Taking $\psi_{\boldsymbol{u}_{h}}=\boldsymbol{F}^{n}_{h}$, $\phi_{\varphi_{h}}=\Phi_{h}^{n}$, $\phi_{B_{h}}=\vartheta_{h}^{n}$ in 24 yields the temporal approximation of 14 as
$\int_{J^{n}}\tau\frac{\partial\mathcal{H}_{h}}{\partial t}=\mathcal{H}_{h}^{n}-\mathcal{H}_{h}^{n-1}=\left(\boldsymbol{u}_{h}^{n}-\boldsymbol{u}_{h}^{n-1},\boldsymbol{F}_{h}^{n}\right)+\left(\varphi_{h}^{n}-\varphi_{h}^{n-1},\Phi_{h}^{n}\right)+\left(B_{h}^{n}-B_{h}^{n-1},\vartheta_{h}^{n}\right)=0,$ (26)
which is zero since skew symmetry is maintained. Proving 24b holds point-wise in space and time yields discrete mass conservation.
Proposition 4.1.
The continuity equation 24b holds point-wise in space and time.
Proof.
Assuming continuity of time, 8b shows $(\partial_{t}\varphi_{h}+\nabla\cdot\boldsymbol{F}_{h})(x,\cdot)\in\mathbb{V}_{2}$ holds point-wise in space. Interpolation of the initial condition yields $\varphi_{h}^{0}\in\mathbb{V}_{2}$. Since $\varphi_{h}^{n},\nabla\cdot\boldsymbol{F}_{h}^{n}\in\mathbb{V}_{2}$ for $n=1,2,\dots,N$, it is clear to see $(\varphi_{h}^{n}-\varphi_{h}^{n-1}+\tau\nabla\cdot\boldsymbol{F}_{h}^{n})\in\mathbb{V}_{2}$ holds point-wise in time. ∎
A key novelty of the fully discrete formulation in 24 is that the diagnostic buoyancy is represented as a continuous linear polynomial in time. Consequently, discrete entropy is a continuous cubic polynomial in time, which is not conserved point-wise in time by the chosen Poisson integrator [Cohen2011]. We now prove the loss in exact entropy conservation depends on the accuracy of the temporal approximation.
Proposition 4.2.
Entropy conservation has a drift of size $\mathcal{O}(\tau)$, that is, $|\mathcal{S}_{h}^{n}-\mathcal{S}_{h}^{0}|=\mathcal{O}(\tau)$.
Proof.
[…]
$\mathcal{S}_{h}^{n}$ $=\int_{\Omega}\frac{1}{2}\left[\left(b_{h}^{n}-b_{h}\right)+b_{h}\right]\left[\left(b_{h}^{n}-b_{h}\right)+b_{h}\right]\left[\left(\varphi_{h}^{n}-\varphi_{h}\right)+\varphi_{h}\right],$ (27a)
$=\int_{\Omega}\frac{1}{2}b_{h}b_{h}\varphi_{h}+\int_{\Omega}\frac{1}{2}\left(b_{h}^{n}-b_{h}\right)\left(b_{h}^{n}-b_{h}\right)\left(\varphi_{h}^{n}-\varphi_{h}\right)+\dots$ (27b)
$=\mathcal{S}_{h}^{0}+\mathcal{O}(\tau).$ (27c)
Thus, $|\mathcal{S}_{h}^{n}-\mathcal{S}_{h}^{0}|=\mathcal{O}(\tau)$ and this completes the proof. ∎
Subsequent to Proposition 4.2, the relation in 17 is not discretely preserved point-wise in time. To demonstrate the importance of 25 in regards to discrete entropy conservation, we consider the temporal approximation of 21 as
$\int_{J^{n}}\tau\frac{\partial\mathcal{S}_{h}}{\partial t}=\mathcal{S}_{h}^{n}-\mathcal{S}_{h}^{n-1}=\frac{1}{2}\left([b_{h}b_{h}]^{n-1/2},\nabla\cdot\boldsymbol{F}_{h}^{n}\right)-\frac{1}{2}\left(\widetilde{b_{h}^{n}}b_{h}^{n-1/2},\nabla\cdot\boldsymbol{F}_{h}^{n}\right)-s_{\mathrm{up}}(\boldsymbol{F}_{h}^{n},b_{h}^{n-1/2},b_{h}^{n-1/2}).$ (28)
Using $\widetilde{\phi}_{h}=\nabla\cdot\boldsymbol{F}_{h}^{n}$ in 25 yields $\partial\mathcal{S}_{h}/\partial t=0$ when $\alpha=0$, and ${\partial\mathcal{S}_{h}}/{\partial t}<0$ when $\alpha>0$. We emphasise the choice of $\widetilde{b_{h}^{n}}$ means no spatial or temporal entropy conservation errors arise in the forcing terms. Thus, the only source of entropy conservation error is the approximation of temporal derivatives. Numerical experiments in Section 5 below support this analysis.
4.4. Constrained formulation
Temporal conservation of Casimirs can be enforced using a global constraint. In the context of discrete entropy conservation, we find $b_{h}^{n}\in\mathbb{V}_{2}$, a minimiser of,
$\mathcal{J}(b_{h}^{n})$ $=\frac{1}{2}\left(b_{h}^{n}\varphi_{h}^{n},b_{h}^{n}\right)-\left(B_{h}^{n},b_{h}^{n}\right)\quad\text{such that }\quad\mathcal{S}_{h}^{n}-\mathcal{S}_{h}^{n-1}=0.$ (29)
Equation 29 can be written as: find $\left[b_{h}^{n},\lambda\right]\in\mathbb{V}_{2}\times\mathbb{R}$ such that
$\left(b_{h}^{n}\varphi_{h}^{n},\phi_{b_{h}}\right)-\left(B_{h}^{n},\phi_{b_{h}}\right)+\lambda\left(b_{h}^{n}\varphi_{h}^{n},\phi_{b_{h}}\right)+\frac{1}{2}\mu\left(b_{h}^{n}\varphi_{h}^{n},b_{h}^{n}\right)$ $=\mu\mathcal{S}_{h}^{n-1},$ $\forall\left[\phi_{b_{h}},\mu\right]\in\mathbb{V}_{2}\times\mathbb{R},$ (30)
where $\lambda\in\mathbb{R}$ is a Lagrange multiplier. Diagnosing $b_{h}^{n}$ using 30 instead of 10d means the analysis of temporal entropy conservation in Section 4.3 is trivial, since we explicitly enforce $\tau\partial\mathcal{S}_{h}/\partial t=\mathcal{S}_{h}^{n}-\mathcal{S}_{h}^{n-1}=0$ point-wise in time. It follows from Proposition 4.2 that $\lambda$ is sufficiently small and corrects an $\mathcal{O}(\tau)$ drift in exact entropy conservation. Further, using 30 perturbs $b_{h}^{n}$ everywhere in 24. Since skew symmetry is maintained independent of $b_{h}^{n}$, discrete energy conservation holds as in 26 by construction of the chosen Poisson integrator [Cohen2011].
4.5. Quasi-Newton approach
The fully discrete system 24 is an implicit non-linear problem, for which using a quasi-Newton approach to approximate the Jacobian can reduce computational expense. The Jacobian is computed once per time step as
$\langle\mathbf{J}(\boldsymbol{\delta u_{h}},\delta\varphi_{h},\delta B_{h}),(\psi_{\boldsymbol{u}_{h}},\phi_{\varphi_{h}},\phi_{B_{h}})\rangle$ $=\left(\boldsymbol{\delta u}_{h},\psi_{\boldsymbol{u}_{h}}\right)+\frac{\tau}{2}\left(\omega_{h}^{n-1},\boldsymbol{\delta u}_{h}^{\perp}\cdot\psi_{\boldsymbol{u}_{h}}\right)$ (31)
$-\frac{\tau}{2}\left(\nabla\cdot\psi_{\boldsymbol{u}_{h}},\frac{\delta B_{h}}{2}\right)-\frac{\tau}{2}\left(\nabla\cdot\psi_{\boldsymbol{u}_{h}},b_{h}^{n-1}\frac{\delta\varphi_{h}}{2}\right)$
$+\left(\delta\varphi_{h},\phi_{\varphi_{h}}\right)+\frac{\tau}{2}\left(\nabla\cdot\boldsymbol{\delta u}_{h},\varphi_{h}^{n-1}\phi_{\varphi_{h}}\right)$
$+\left(\delta B_{h},\phi_{B_{h}}\right)+\frac{\tau}{2}\left(\nabla\cdot\boldsymbol{\delta u}_{h},b_{h}^{n-1}\varphi_{h}^{n-1}\phi_{B_{h}}\right),$
where $\omega_{h}^{n-1}\in\mathbb{X}_{01}$ is diagnosed by setting $\omega_{h}^{n-1}=q_{h}^{n-1}\varphi_{h}^{n-1}$ in 10e. Equation 31 originates from an implicit midpoint discretisation of the linear thermal shallow water equations, in which fast linear waves are accounted for and non-linear coupling terms are omitted [Eldred2019, Lee2022, BauerCotter2018, Wimmer2020]. The linearisation in the current study differs from other works [Eldred2019, Lee2022] as we consider perturbations from the solution at $t^{n-1}$ by using $\varphi_{h}^{n-1},b_{h}^{n-1}$ in 31 instead of the initial usual mean flow state. In doing so, we account for turbulent states which significantly deviate from the resting steady state as simulations progress. Including upwinded numerical fluxes may result in solver convergence issues when simulating a mature turbulent state (see Section 5.2). In these situations, the Jacobian is recomputed at every non-linear iteration and includes approximated centred and upwinded fluxes.
4.6. Implementation remarks
The quasi-Newton approach is implemented using Gridap.jl [Badia2020, Verdugo2022], a Julia package for finite element computations. The implicit time stepping framework within Gridap.jl is modified to solve the linear diagnostic problem prior to evaluating the residual of the non-linear prognostic problem [Eldred2019]. The linear diagnostic problem is constructed as explicit. The current guesses for $\boldsymbol{u}_{h}^{n},\varphi_{h}^{n},B_{h}^{n}$ are used to:
  1. (1)
    Evaluate $\boldsymbol{F}^{n}_{h}\in\mathbb{V}_{1}$, $\Phi^{n}_{h},\vartheta^{n}_{h},b_{h}^{n}\in\mathbb{V}_{2}$, $q_{h}^{n-1/2}\in\mathbb{V}_{0}$ exactly in time via 10.
  2. (2)
    Construct $b_{h}^{n-1/2}$ and $[b_{h}b_{h}]^{n-1/2}$, and diagnose $\widetilde{b_{h}^{n}}$ via 25.
We use GridapDistributed.jl [Badia2022], GridapPETSc.jl [GridapPETSc] and GridapSolvers.jl [GridapSolvers], to conduct tests on the Gadi@NCI Australian supercomputer. The source code for this study is available on Zenodo [Zenodo].
Like other upwinded schemes [Brezzi2004], the upwinding parameter $\alpha$ depends on the direction of flow, and requires a numerical implementation of the signum function, $\operatorname{sign}(x)$. One possibility is
$\operatorname{sign}(x)=\begin{cases}1&x>\epsilon,\\ -1&x<-\epsilon,\\ 0&\text{otherwise},\end{cases}$ (32)
which converges to $\operatorname{sign}(x)$ as $\epsilon\rightarrow 0$ for $\epsilon\geq 0$. However 32 is a non-differentiable function that may cause issues in numerical solvers [Badia2017]. An alternative approach is to approximate $\operatorname{sign}(x)$ by a smooth function [Badia2017] such as
$\operatorname{sign}(x)=\frac{x}{\sqrt{x^{2}+\epsilon^{2}}},$ (33)
which also converges to $\operatorname{sign}(x)$ as $\epsilon\rightarrow 0$. We refer to 32 and 33 as the hard and soft signum functions respectively, and explore their effects in the numerical experiments below. The value of $\epsilon$ is based on values of $||\boldsymbol{F}_{h}^{n}\cdot\boldsymbol{n}^{+}||$ from experiments using the centred scheme.
5. Numerical experiments
To examine the ability of the new finite element approximation to conserve system invariants over long periods of time, we first evaluate convergence under $h$-$p$ refinement and then consider transient test cases that exhibit turbulent flows.
5.1. Convergence test
Convergence is assessed using a steady zonal thermogeostrophic balance test [Eldred2019]. In all simulations, the time step is $\tau=Ch$ where $C=\mathrm{CFL}/(p^{2}\sqrt{g\varphi_{0}})$, $p$ is the spatial polynomial order, $h$ is the spatial step, and $\sqrt{g\varphi_{0}}$ represents the mean speed of gravity waves for which pressure gradient forces dominate [Vallis2006]. Simulations are conducted in a reference frame where the domain length is an order 1 quantity. This is generated by rescaling the initial condition using the length scale $L_{0}$, time scale $T_{0}$, velocity scale $U_{0}$, depth scale $H_{0}$, and buoyancy scale $b_{0}$. This yields the Rossby number $\mathrm{Ro}=U_{0}/(T_{0}L_{0})$ and the Burgers number $\mathrm{Bu}=b_{0}H_{0}/(T_{0}^{2}L_{0}^{2})$.
The rescaled initial condition is
${u}_{1}$ $=\cos({x}_{2}),$ ${u}_{2}$ $=0,$ ${\varphi}$ $=1-\frac{\mathrm{Ro}}{\mathrm{Bu}}\sin({x}_{2}),$ ${b}$ $=1+c{\varphi},$ (34)
where ${\boldsymbol{u}}=(u_{1},u_{2})$, ${B}={b}{\varphi}$, ${x}_{1,2}$ are the horizontal and vertical spatial coordinates respectively, and $c=0.05$. In the rescaled frame, ${\Omega}=[0,2\pi]^{2}$ is doubly periodic, ${g}=\varphi_{0}=1$, ${f}=\mathrm{Ro}/\mathrm{Bu}$, and the duration corresponds to 5 days [Ricardo2024-dg]. The scaling parameters are: $L_{0}=6371120\text{ m}$, $T_{0}=6.147\times 10^{-5}\text{ s}^{-1}$, $U_{0}=20\text{ m s}^{-1}$, $H_{0}=5960\text{ m}$, and $b_{0}=9.80616\text{ m s}^{-1}$.
Convergence of the centred scheme is tested using the relative $L^{2}$ norm between the initial and final solutions in the absence of stabilisation [Eldred2019]. Fig. 1(a) shows the expected third and fourth order convergence for a range of spatial and temporal resolutions. For each line in Fig. 1, ${h}$ and ${\tau}$ reduce simultaneously at rate $C=\mathrm{CFL}/p^{2}$, where $\mathrm{CFL}=0.2$ for $p=1$ and $\mathrm{CFL}=0.1$ for $p=2$. Similar convergence rates are observed in Fig. 1(b) for the upwinded scheme.
Figure 1. Convergence of the zonal thermogeostrophic test case showing relative $L^{2}$ error between initial and final solutions for (a) the centred scheme and (b) the upwinded scheme (hard signum, $\epsilon=10^{-4}$), with 16, 32, 64, 128 spatial elements in each case.
5.2. Case study 1: thermal instability
We now consider transient case studies to confirm adequate conservation of system invariants under turbulent flow conditions and over long time periods relative to the onset of geostrophic turbulence. Discrete conservation is demonstrated using normalised values of discrete energy 13, entropy 16 and mass ($\mathcal{M}_{h}(t)=\int_{\Omega}\varphi_{h}$) on a logarithmic scale over time. The change in entropy due to forcing terms, $\partial\mathcal{S}_{h}/\partial t$ in 28, is also shown. In all simulations, the tolerance of the non-linear solver is $10^{-12}$, and conservation errors up to $\mathcal{O}(10^{-12})$ are reasonable.
The first transient case study involves a thermal instability [Eldred2019, Gouzien2017, Zeitlin2018, Kurganov2021]. Other studies simulate this test case using an upwinded finite volume scheme [Kurganov2021, Gouzien2017, Zeitlin2018]. A previous finite element study shows energy conservation at early times [Eldred2019], but entropy conservation, the evolution of complex non-linear flows, and the effect of upwinding is not considered. We extend these works using the novel centred and upwinded fluxes to simulate the thermal instability test case a over long time period while examining conservation of system invariants.
The initial condition is
${u}_{1}$ $=-U_{0}r\exp\left(\frac{1-r^{\beta}}{\beta}\right)\sin(\phi)+\varepsilon,$ ${u}_{2}$ $=U_{0}r\exp\left(\frac{1-r^{\beta}}{\beta}\right)\cos(\phi)+\varepsilon,$ (35)
${\varphi}$ $=1-\varepsilon,$ ${b}$ $=1-2\frac{\mathrm{Ro}}{\mathrm{Bu}}\left(\exp\left(\frac{1-r^{2}}{2}\right)+\frac{\mathrm{Ro}}{2}\exp\left(1-r^{2}\right)\right)+\varepsilon,$
where
$\varepsilon$ $=0.01\exp\left(-60(r-r_{c})^{2}\right)\sin\left(6\pi(r-r_{c})\right)\cos(4\phi),$ (36)
is a perturbation, ${r}^{2}={x}_{1}^{2}+{x}_{2}^{2}$ is the radial distance, $\tan(\phi)=x_{2}/x_{1}$ is the polar angle, and $\Omega=[-4,4]^{2}$ is doubly periodic. The duration, $T=100$, is sufficiently long for a mature turbulent state to develop. The parameters are: $\beta=2$, $r_{c}=0.5$, $g=\varphi_{0}=f=L_{0}=H_{0}=b_{0}=T_{0}=\mathrm{Bu}=1$, and $U_{0}=\mathrm{Ro}=0.1$.
Fig. 2 shows the evolution of the thermal instability for the centred and upwinded schemes. Similar to other studies that also consider this test case [Eldred2019, Gouzien2017], the initial growth evident in Fig. 2(a)–(d) reflects the wave number of 4 used in 36, and non-linear saturation of the instability occurs as the simulation progresses (Fig. 2(e)–(h)). Comparing the buoyancy field at $t=25$ (Fig. 2(a,b)) that obtained at $t=100$ (Fig. 2(g,h)) shows a large variation in buoyancy. Regardless, both the centred and upwinded schemes resolve the highly non-linear flow that develops as the simulation progresses. This is a direct consequence of the new linearised Jacobian 31, which allows for robust convergence for both centred and upwinded fluxes in the presence of significant variations from the initial mean flow state. As expected, including upwinded fluxes (Fig. 2(b,d,f,g)) suppresses spurious oscillations which arise in the absence of stabilisation (Fig. 2(a,c,e,g)).
We now assess the ability of the centred and upwinded schemes to conserve system invariants in the presence of turbulent flow. In the absence of stabilisation, Fig. 3(a,b) shows energy and mass conservation, while Fig. 3(c) demonstrates the entropy conservation error increases to $\mathcal{O}(10^{-8})$ over time even though ${\partial\mathcal{S}_{h}}/{\partial t}$ is machine zero for the entire duration (Fig. 3(d)). This supports the analysis in Section 4.3 above that entropy is not conserved point-wise in time. While an entropy conservation error of $\mathcal{O}(10^{-8})$ may seem large relative to energy and mass, it is comparable with other studies that consider temporal conservation of quadratic tracer invariants [Lee2022]. Thus, we conclude the centred scheme stably resolves turbulent dynamics over long periods of time and conserves system invariants.
For the upwinded scheme, the spikes in energy and mass evident in Fig. 3(a,b) correspond to instances when the non-linear solver fails to converge. Values of $\epsilon=10^{-3}$ and $\epsilon=10^{-4}$ are chosen to approximate $\operatorname{sign}(x)$ in the the upwinded scheme since $||\boldsymbol{F}_{h}^{n}\cdot\boldsymbol{n}^{+}||\approx 0.1$ for the centred scheme. When using the hard signum function with $\epsilon=10^{-4}$, solver convergence issues arise at $t\approx 45$. In contrast, using the soft signum function for $\epsilon=10^{-4}$ and $\epsilon=10^{-3}$ delays the onset of solver convergence issues to $t\approx 70$ and $t\approx 98$ respectively. Therefore, using the soft signum function and increasing $\epsilon$ assists solver convergence, in particular conservation of energy and mass, as turbulent flow dynamics evolve.
A consequence of turbulent flow dynamics is that the direction of flow may change within discrete time intervals. Since the Jacobian is recomputed at every non-linear iteration and includes both approximated centred and upwinded fluxes, we make every effort to minimise solver convergence issues regardless of the choice of signum function or $\epsilon$ value. As with any numerical simulation of partial differential equations, using an increasingly fine computational mesh and sophisticated non-linear solver can reduce convergence issues. We acknowledge non-linear solver issues may arise given the exceedingly small numbers required to demonstrate conservation in the context of highly turbulent flow.
The logarithmic axis in Fig. 3(c,d) helps exemplify multiple orders of magnitude difference in entropy conservation between the centred and upwinded schemes. Using a linear vertical axis to represent conservation error shows that indeed ${\partial\mathcal{S}_{h}}/{\partial t}<0$ for the upwinded scheme, and entropy is monotonically damped (not shown). This supports the analysis of temporal entropy conservation in Section 4.3.
Figure 2. Evolution of buoyancy $b_{h}$ for the thermal instability case study where (a,c,e,g) relate to the centred scheme and (b,d,f,h) correspond to the upwinded scheme (hard signum, $\epsilon=10^{-4}$). Snapshots are shown at $t=25,50,75,100$ where time is unit-less. Parameters: $n_{s}=192$ spatial elements, $p=1$ spatial finite elements, $\mathrm{CFL}=0.2$.
Figure 3. Conservation errors for the thermal instability case study. (a,b,c) show normalised values of energy, mass and entropy, and (d) illustrates the change in entropy due to forcing terms 21. Each figure compares the centred and upwinded schemes for different approximations of signum and $\epsilon$ values. The vertical axis is logarithmic and the horizontal axis shows unit-less time. Parameters: $n_{s}=192$ spatial elements, $p=1$ spatial finite elements, $\mathrm{CFL}=0.2$.
5.3. Case study 2: double vortex
The second transient case study describes a double vortex [Eldred2019]. Test cases involving vortex pair interactions are often used to assess numerical schemes for the rotating shallow water equations [Giorgetta2009, Reich2006, Mcrae2013, Staniforth2006]. A previous study adapted such a test case to the thermal shallow water equations by considering the time evolution of two vortices interacting in a zonally varying buoyancy field [Eldred2019]. While this previous study considered an energy conserving finite element formulation [Eldred2019], entropy conservation was not of concern. Thus, we use the novel centred fluxes proposed in the current study to achieve semi-discrete entropy conservation, and demonstrate the use of Lagrange multipliers to rectify small losses in exact entropy conservation.
The rescaled initial condition is
${u}_{1}$ $=-\eta_{21}\varepsilon_{1}-\eta_{22}\varepsilon_{2},$ ${u}_{2}$ $=\eta_{11}\varepsilon_{1}+\eta_{12}\varepsilon_{2},$ (37)
${\varphi}$ $=1-\varphi_{c}\left(\varepsilon_{1}+\varepsilon_{2}-4\pi\sigma^{2}\right),$ ${b}$ $=1+c\sin\left(2\pi{x}_{1}-\pi\right),$
where
$\varepsilon_{j}$ $=\exp\left(-\frac{1}{2}(\gamma_{1j}^{2}+\gamma_{2j}^{2})\right),$ $\gamma_{ij}$ $=\frac{1}{\pi\sigma}\sin\left(\pi({x}_{i}-{c}_{j})\right),$ $\eta_{ij}$ $=\frac{1}{2\pi\sigma}\sin\left(2\pi({x}_{i}-{c}_{j})\right),$ (38)
for $i,j=1,2$, $\varphi_{c}=h_{0}/H_{0}$, $\sigma=3/40$, ${c}_{1}=0.4$, ${c}_{2}=0.6$ and $c=0.05$ [Eldred2019]. In the rescaled frame, ${\Omega}=[0,1]^{2}$ is doubly periodic, ${g}={\varphi}_{0}=1$, and ${f}$ is computed such that $\mathrm{Bu}$ is constant between the rescaled and unscaled frames. The duration is chosen to allow mature turbulent structures to develop. The scaling parameters are: $L_{0}=5\times 10^{6}$ m, $T_{0}=6.147\times 10^{-5}$ s-1, $h_{0}=75\text{ m}$, $H_{0}=750\text{ m}$, $b_{0}=9.80616\text{ m s}^{-2}$ and $U_{0}=(b_{0}h_{0})/(T_{0}L_{0}\sigma)$ [Eldred2019].
Fig. 4 shows the centred scheme adequately resolves turbulent flow structures over long periods of time with minimal grid scale noise. Similar to other studies that consider vortex pair interactions [Giorgetta2009, Reich2006, Mcrae2013, Staniforth2006, Eldred2019], we observe the formation of two distinct vortices (Fig. 4(a)) which collide as a mature turbulent state is reached (Fig. 4(f)). The ability of the centred scheme to resolve such complex features is further evidence that the new linearised Jacobian 31 provides robust convergence with respect to large variations from the initial mean flow.
Figure 4. Evolution of buoyancy $b_{h}$ for the double vortex case study, simulated with the centred scheme. Snapshots are shown at $t=0.75,1.5,2,3,4,5$ where $t$ is unit-less. Parameters: $n_{s}=192$ spatial elements, $p=1$ spatial finite elements, $\mathrm{CFL}=0.2$.
For the centred scheme, Fig. 5(a,b) shows conservation of energy and mass to machine precision. Similar to Section 5.2, Fig. 5(c) shows a drift in entropy conservation from machine precision to $\mathcal{O}(10^{-8})$ as the simulation progresses even though $\partial\mathcal{S}_{h}/\partial t$ is machine zero (Fig. 5(d)). Thus, we conclude entropy is not conserved point-wise in time.
Figure 5. Conservation errors for the double vortex case study. (a,b,c) show normalised values of energy, mass and entropy, and (d) illustrates the change in entropy due to forcing terms 21. Each figure compares the centred (blue) and constrained schemes (orange). The vertical axis is logarithmic and the horizontal axis shows unit-less time. Parameters: $n_{s}=64$ spatial elements, $p=1$ spatial finite elements, $\mathrm{CFL}=0.2$.
To achieve exact entropy conservation, we use the constrained formulation in 30, for which entropy conservation is explicitly enforced using a Lagrange multiplier. Conservation errors in energy, mass and $\partial\mathcal{S}_{h}/\partial t$ are are comparable between the centred and constrained formulations (Fig. 5(a,b,d)), and Fig. 5(c) shows entropy is conserved within the tolerance of the non-linear solver. In situations where exact entropy conservation is required, we advise use of the constrained formulation.
6. Conclusion and future work
In this study, we develop a new finite element discretisation for the thermal shallow water equations. Novel centred and upwinded fluxes are developed to include a specific formulation of buoyancy, $\widetilde{b_{h}}$, at the semi-discrete level. We prove the centred fluxes are energy and entropy conserving, while the upwinded fluxes are energy conserving and entropy damping. Using a Poisson time integrator [Cohen2011] and an original temporal representation of discrete thermodynamic quantities, we construct a fully discrete system that conserves entropy up to the accuracy of the temporal approximation, and conserves energy by construction. In particular, the temporal formulation of $\widetilde{b_{h}}$ ensures the forcing terms conserve entropy in time, such that temporal conservation errors only arise from approximating time derivatives. The developed quasi-Newton approach utilises the solution at progressive time levels instead of the typical mean flow state [Eldred2019, Lee2022] to capture the evolution of turbulent dynamics. Transient case studies involving a thermal instability [Eldred2019, Gouzien2017, Zeitlin2018, Kurganov2021] and double vortex [Eldred2019, Giorgetta2009] illustrate the ability of the novel centred and upwinded fluxes, and linearised Jacobian to stably resolve complex flow structures over long periods of time.
A consequence of the chosen Poisson time integrator [Cohen2011] is that cubic Casimirs are not preserved for non-canonical Hamiltonian systems. Thus, discrete entropy is not conserved point-wise in time. To address this, we propose a constrained formulation using Lagrange multipliers which numerical experiments confirm corrects small temporal losses in exact entropy conservation. While Poisson time integrators that preserve higher order Casimirs for canonical Hamiltonian systems have recently been developed [Modin2020], the extension to non-canonical Hamiltonian systems is an open problem.
There are many potential avenues to extend the framework presented in this study. We take a fundamental approach and showcase the new finite element scheme using planar test cases. One avenue for extension is to simulate spherical test cases [Ricardo2024-dg]. This is feasible since the numerical method and analysis presented in this study does not depend on the choice of coordinate system. The quasi-Newton approach and linearised Jacobian developed in this study could be extended to other problems that consider quasi-Newton methods in conjunction with multigrid preconditioners for atmospheric models [Maynard2020]. We always consider skew symmetric formulations of the thermal shallow water equations. So the developed numerical schemes and analysis could be interpreted in the context of the compressible Euler equations [Ricardo2023]. This would require careful consideration of the spatial and temporal approximation of relevant thermodynamic quantities and variational derivatives, and is left for future consideration.
Acknowledgements
This research is supported by the Commonwealth of Australia as represented by the Defence Science and Technology Group of the Department of Defence. This research is also funded by the Australian Government through the Australian Research Council (project numbers DP210103092 and DP220103160). This work is supported by computational resources provided by the Australian Government through NCI under the National Computational Merit Allocation Scheme (NCMAS), ANU Merit Allocation Scheme, and the Monash-NCI partnership program.
Author contribution
All authors contributed equally to the design of the study. Tamara A. Tambyah performed numerical simulations and drafted the article. All authors gave approval for publication.
Declaration of competing interest
The authors declare they have no competing interests.
Data availability
This articles does not contain any additional data. The source code for this study is available on Zenodo [Zenodo].
[…]